OpenStudy (anonymous):

Choose the correct simplification of (x9)2. x11 x7 x81 x18

OpenStudy (goformit100):

wHAT IS THE QUESTION?

hartnn (hartnn):

\(\huge (x^m)^n= x^{mn}\)

OpenStudy (phi):

\[(x^9)^2\] this means \[x^9 \cdot x^9\] x^9 means x*x*x* x*x*x* x*x*x and x^9 * x^9 = x*x*x* x*x*x* x*x*x* x*x*x* x*x*x* x*x*x count the x's

OpenStudy (phi):

of course, if you do that a few times, you will figure out that 2*9=18 is the faster way